Binomial Distribution

In a survey, 25% of the people interviewed said that they bought their refrigerator during the last six months. If eleven people are selected at random, find the probability that exactly six of these people bought their refrigerator during the last six months.


1
Expert's answer
2021-05-07T09:19:02-0400

Let x be a random variable denoting the number of people who bought a refrigerator during the last 6 months.

p=0.25; q=1-p=1-0.25=0.75

Since there are two events, apply the binomial distribution.

"P(X=x) ={n \choose x} p^x q^{n-x}"

When n=11; p(X=6) will be ;

"P(X=6)={11 \choose 6}×0.25^6×0.75^{11-6}"

=0.0268